# Gauge Symmetries, Contact Reduction, and Singular Field Theories

### v1

@@ XIII Conclusions and Outlook
Many of the most mathematically-rich models in contemporary theoretical physics possess, as a central component, gauge symmetries; such theories are described by singular Lagrangians, and over the course of this article, we have provided a basic introduction to gauge theories, from the perspective of geometric mechanics. For systems of particles, the appropriate formalism was that of pre-symplectic geometry, while a field-theoretic discussion required the introduction of the fibred manifolds of pre-multisymplectic geometry. We have demonstrated that singular theories which exhibit invariance under global scaling transformations may be subjected to a reduction process similar to that which has been studied for regular systems, provided that degeneracies are controlled, through use of a suitable constraint algorithm. We observed that this constraint procedure could be implemented pre- or post-reduction; from a physical perspective, this is to be expected: if the processes of contact reduction and phase space restriction were non-commutative, this would imply that the constraint functions had, in some sense, information about the global scaling variable. Had the reduction of a constrained system led to a theory distinct from that obtained via the constraint of a reduced theory, then any system described by a degenerate Lagrangian would offer a means to detect changes in global scale, which is a highly unphysical outcome.
@@
Having described the reduction of singular field theories, we then provided an example of an effective non-Abelian gauge theory coupled to a dilaton-like field; it was observed that this real scalar field was in fact redundant, and could therefore be excised. The elimination of this degree of freedom was, at first sight, relatively alarming, particularly because the coupling parameters of a theory are often fixed by the expectation value of such scalar fields. However, we argued that the strength of an interaction is something that must necessarily be deduced via empirical methods; as such, any hypothetical experiment, conducted with the aim of measuring such parameters will be insensitive to a rescaling of their values, since the apparatus itself will suffer an identical change in scale.
@@
The results presented throughout constitute a framework within which to analyse singular theories; it is known that, upon making a conformal decomposition of the spacetime metric, the Einstein-Hilbert Lagrangian possesses a scaling symmetry, corresponding precisely to the conformal factor [34]. To our knowledge, this scaling symmetry has not been studied within the first-order Palatini formalism, for which the multisymplectic Lagrangian description has been developed in [99] and [100]. From the results of the present work, we know that it is possible to implement the geometrical constraint algorithm after having eliminated the scaling degree of freedom. As such, it will be of great interest to examine any simplifications that arise, as a result of working with a reduced ontology, in which we retain only those degrees of freedom required to close the algebra of dynamical observables.
In general, we distinguish between two principal approaches to the treatment of classical field theory. On the one hand, as is customary when attempting to construct a quantum analogue of the theory, one may consider the temporal evolution of the initial data, defined on a particular choice of Cauchy slice [101]. This ‘canonical’ formulation studies spatially-dependent fields, defined at fixed instants of time, and has the notably inconvenient features of possessing an infinite-dimensional phase space, and sacrificing manifest Lorentz covariance [102]. The ADM formalism of General Relativity is a constrained Hamiltonian theory that adopts just such an approach, considering the (globally-hyperbolic) spacetime manifold to be foliated by spacelike hypersurfaces [103, 104].
The Hamilton De-Donder Weyl formalism we have employed throughout makes no distinction between the spacetime coordinates, which immediately guarantees the Lorentz covariance that was absent in the canonical formulation. Having established a framework within which to treat constrained theories, it will be highly illustrative to contrast the constraint of a contact-reduced action for General Relativity with the more familiar canonical approach.

### v2

@@ XIII Conclusions and Outlook
Many of the most mathematically-rich models in contemporary theoretical physics possess, as a central component, gauge symmetries; such theories are described by singular Lagrangians, and over the course of this article, we have provided a basic introduction to gauge theories, from the perspective of geometric mechanics. We have demonstrated that singular theories which exhibit invariance under global scaling transformations may be subjected to a reduction process similar to that which has been studied for regular systems. Degeneracies must be controlled, and we found that the constraint procedure required to do so could be implemented pre- or post-reduction. From a physical perspective, this is unsurprising: had the processes of contact reduction and phase space restriction been non-commutative, this would imply that the constraint functions had, in some sense, information about the global scaling variable. If the reduction of a constrained system led to a theory distinct from that obtained via the constraint of a reduced theory, then any system described by a degenerate Lagrangian would offer a means to detect changes in global scale, which is a highly unphysical outcome.
@@
Having described the reduction of singular field theories, we provided an example of an effective non-Abelian gauge theory coupled to a dilaton-like field; it was observed that this real scalar field was in fact redundant, and could therefore be excised. The elimination of this degree of freedom was, at first sight, relatively alarming, particularly because the coupling parameters of a theory are often fixed by the expectation values of such scalar fields. However, we argued that the strength of an interaction is something that is necessarily deduced via empirical methods; as such, any hypothetical experiment, conducted with the aim of measuring such parameters will be insensitive to a rescaling of their values, since the apparatus itself will suffer an identical change in scale.
@@
The results presented throughout constitute a framework within which to analyse singular theories; it is known that, upon making a conformal decomposition of the spacetime metric, the Einstein-Hilbert Lagrangian possesses a scaling symmetry, corresponding precisely to the conformal factor [34]. To our knowledge, this scaling symmetry has not been studied within the first-order Palatini formalism, for which the multisymplectic Lagrangian description has been developed in [89] and [90]. From the results of the present work, we know that it is possible to implement the geometrical constraint algorithm after having eliminated the scaling degree of freedom. It will be of great interest, therefore, to examine any simplifications or novelties that arise as a result of working with a reduced ontology, particularly at those points where the standard Einstein-Hilbert action provides ill-defined solutions, where scales become singular. Containing no reference to such scales, our contact-reduced action will not suffer the same pathologies, and will therefore provide an arena in which to explore these singular points.
Finally, more work is required to ascertain how one should construct a field-theoretic generalisation of the promotion of coupling parameters to dynamical variables, discussed at the end of section (XII). Having found such a prescription, it will, in principle, be possible to analyse a much broader class of gauge theories, such as the Standard Model.